### v1

# A congruence subgroup property for symmetric mapping class groups

Abstract.
We prove the congruence subgroup property for the centralizer of a finite subgroup $G$ in the mapping class group of a hyperbolic oriented and connected surface of finite topological type $S$ such that the genus of the quotient surface $S/G$ is at most $2$. As an application, we show that torsion elements in the mapping class group of a surface of genus $\leq 2$ are conjugacy distinguished.
AMS Math Classification: 57K20, 20F65, 14G32.
1. Introduction
Let $S$ be an oriented connected surface of negative Euler characteristic and finite type and let $G$ be a finite group acting faithfully on $S$. The group $G$ then identifies with a finite subgroup of the mapping class group ${\Gamma}(S)$ of the surface $S$ and let us denote by ${\Gamma}_{G}(S)$ the centralizer of $G$ in ${\Gamma}(S)$. Let $\Pi$ be the fundamental group of $S$ for some choice of base point. There is a natural faithful representation $\rho_{S}\colon\thinspace{\Gamma}(S)\hookrightarrow\operatorname{Out}(\Pi)$ which restricts to a faithful representation $\rho_{S,G}\colon\thinspace{\Gamma}_{G}(S)\hookrightarrow\operatorname{Out}(\Pi)$.
For an abstract group $L$, let us denote by $\widehat{L}$ its profinite completion. The homomorphisms $\rho_{S}$ and $\rho_{S,G}$ then induce the homomorphisms of profinite groups:
(1) $\hat{\rho}_{S}\colon\thinspace{\widehat{\Gamma}}(S)\to\operatorname{Out}({\widehat{\Pi}})\hskip 28.45274pt\mbox{and}\hskip 28.45274pt\hat{\rho}_{S,G}\colon\thinspace{\widehat{\Gamma}}_{G}(S)\to\operatorname{Out}({\widehat{\Pi}}).$
The congruence subgroup problem for the mapping class group ${\Gamma}(S)$ (resp. for ${\Gamma}_{G}(S)$) asks whether the homomorphism $\hat{\rho}_{S}$ (resp. $\hat{\rho}_{S,G}$) is faithful. The first question is known to have a positive answer for $g(S)\leq 2$ (cf. [1], [2], [6]), where $g(S)$ denotes the genus of $S$.
Let $S_{/G}:=S/G$ be the quotient of the surface $S$ for the action of $G$ and let $p_{G}\colon\thinspace S\to S_{/G}$ be the quotient map. Let $R$ and $B$ be, respectively, the ramification and the branch locus of $p_{G}$ and put $S^{\circ}:=S\smallsetminus R$ and $S_{/G}^{\circ}:=S_{/G}\smallsetminus B$. The main result of the paper is then (cf. Theorem 4.1 for a slightly more general result):
Theorem A.
Let $S$ be an oriented connected hyperbolic surface of finite type and $G$ a finite group acting faithfully on $S$. If the congruence subgroup property holds for the mapping class group ${\Gamma}(S_{/G}^{\circ})$ of the surface $S_{/G}^{\circ}$, then it holds for the centralizer ${\Gamma}_{G}(S)$ of $G$ in the mapping class group ${\Gamma}(S)$ of $S$.
In particular, we have:
Corollary B.
If $g(S_{/G})\leq 2$, then the congruence subgroup property holds for ${\Gamma}_{G}(S)$.
For $S$ a surface as above, let us denote by $\partial S$ its boundary and by ${\Gamma}(S,\partial S)$ the relative mapping class group of the surface with boundary $(S,\partial S)$, that is to say the group of relative, with respect to the boundary $\partial S$, isotopy classes of orientation preserving diffeomorphisms of $S$. Of course, if $\partial S=\emptyset$, then ${\Gamma}(S,\partial S)={\Gamma}(S)$.
For a subgroup $H$ of a group $G$, we denote by $Z_{H}(G)$ and $N_{H}(G)$, respectively, the centralizer and the normalizer of $H$ in $G$. For a subset $A$ of a profinite group $R$, we denote by $\overline{A}$ the closure of $A$ for the profinite topology of $R$. We then have:
Theorem C.
Let $S$ be an oriented connected hyperbolic surface of finite type of genus $\leq 2$.
  1. (i)
    The natural monomorphism ${\Gamma}(S,\partial S)\hookrightarrow{\widehat{\Gamma}}(S,\partial S)$ induces a bijection between conjugacy classes of solvable finite subgroups. In particular, ${\Gamma}(S,\partial S)$ is solvable finite subgroup conjugacy separable.
  2. (ii)
    Torsion elements are conjugacy distinguished.
  3. (iii)
    For every finite subgroup $G$ of ${\Gamma}(S,\partial S)$, there holds $Z_{{\widehat{\Gamma}}(S,\partial S)}(G)=\overline{Z_{{\Gamma}(S,\partial S)}(G)}$.
    If, moreover, $G$ is solvable, we also have $N_{{\widehat{\Gamma}}(S,\partial S)}(G)=\overline{N_{{\Gamma}(S,\partial S)}(G)}$.
2. Preliminary results
2.1. Normalizers and centralizers of finite subgroups of the mapping class group of a surface
It is clearly not restrictive to assume that $S$ has empty boundary and then that it is obtained from a closed surface removing a finite set of points. With the notations of Section 1, the cover $p_{G}\colon\thinspace S\to S_{/G}$ determines a normal finite index subgroup $K\cong\pi_{1}(S^{\circ})$ of the fundamental group $\Pi_{/G}:=\pi_{1}(S_{/G}^{\circ})$ (we omit base points).
Let ${\Gamma}(S_{/G}^{\circ})_{K}$ be the finite index subgroup of the mapping class group ${\Gamma}(S_{/G}^{\circ})$ which consists of the mapping classes which preserve the normal subgroup $K$ of $\Pi_{/G}$. The normalizer $N_{{\Gamma}(S)}(G)$ of the finite subgroup $G\cong\left.\Pi_{/G}\right/K$ in the mapping class group ${\Gamma}(S)$ is then described by the short exact sequence:
(2) $1\to G\to N_{{\Gamma}(S)}(G)\to{\Gamma}(S_{/G}^{\circ})_{K}\to 1$
Let $Z(G)$ be the center of the group $G$ and let ${\Gamma}^{K}$ be the kernel of the natural representation ${\Gamma}(S_{/G}^{\circ})_{K}\to\operatorname{Out}(\left.\Pi_{/G}\right/K)$. The centralizer ${\Gamma}_{G}(S)\leq N_{{\Gamma}(S)}(G)$ of the subgroup $G$ in the mapping class group ${\Gamma}(S)$ is then described by the short exact sequence:
(3) $1\to Z(G)\to{\Gamma}_{G}(S)\to{\Gamma}^{K}\to 1$
The group $G$ is the covering transformation group of the unramified cover $p_{G}^{\circ}\colon\thinspace S^{\circ}\to S_{/G}^{\circ}$ and thus also identifies with a finite subgroup of the mapping class group ${\Gamma}(S^{\circ})$ of $S^{\circ}$. From the short exact sequences (2) and (3), it follows:
Proposition 2.1.
The natural epimorphism of mapping class groups $q^{\circ}\colon\thinspace{\Gamma}(S^{\circ})\to{\Gamma}(S)$, associated to the embedding of $S^{\circ}$ in $S$, induces the isomorphisms:
$N_{{\Gamma}(S^{\circ})}(G)\cong N_{{\Gamma}(S)}(G)\hskip 28.45274pt\mbox{and}\hskip 28.45274pt{\Gamma}_{G}(S^{\circ})\cong{\Gamma}_{G}(S).$
Proof.
Let us just prove the first isomorphism, the second one is proved similarly. By the short exact sequence (2), applied to the surfaces $S^{\circ}$ and $S$, respectively, there is a commutative diagram with exact rows:
$\begin{array}[]{ccccccc}1\to&G&\to&N_{{\Gamma}(S^{\circ})}(G)&\to&{\Gamma}(S_{/G}^{\circ})_{K}&\to 1\\ &\parallel&&\downarrow^{q^{\circ}}&&\parallel&\\ 1\to&G&\to&N_{{\Gamma}(S)}(G)&\to&{\Gamma}(S_{/G}^{\circ})_{K}&\to 1,\end{array}$
from which the conclusion follows. ∎
2.2. The congruence topology
The congruence topology on ${\Gamma}(S)$ is the subspace topology induced by the standard profinite topology on $\operatorname{Out}({\widehat{\Pi}})$ via the natural monomorphism ${\Gamma}(S)\hookrightarrow\operatorname{Out}({\widehat{\Pi}})$. The congruence topology on ${\Gamma}(S)$ then induces a topology on its subgroups which we also call the congruence topology.
Let ${\Gamma}_{G}^{(m)}(S)$ be the kernel of the natural representation ${\Gamma}_{G}(S)\to\operatorname{GL}(H_{1}(S,\mathbb{Z}/m))$ and let ${\Gamma}^{K,(m)}$ be its image via the natural epimorphism ${\Gamma}_{G}(S)\to{\Gamma}^{K}$. For $m\geq 3$, we have ${\Gamma}_{G}^{(m)}(S)\cap G=\{1\}$ and so the map ${\Gamma}_{G}(S)\to{\Gamma}^{K}$ induces an isomorphism ${\Gamma}_{G}^{(m)}(S)\cong{\Gamma}^{K,(m)}$.
Lemma 2.2.
For all $m\geq 2$, the finite index subgroups ${\Gamma}_{G}^{(m)}(S)$ of ${\Gamma}_{G}(S)$ and ${\Gamma}^{K,(m)}$ of ${\Gamma}(S_{/G}^{\circ})$ are both open for the respective congruence topologies.
Proof.
The statement is obvious for ${\Gamma}_{G}^{(m)}(S)$. For ${\Gamma}^{K,(m)}$, it can be proved with the same argument of [3, Corollary 2.3]. ∎
2.3. The centralizer of a finite subgroup in the profinite completion of a good group
Let $\iota\colon\thinspace G\to\widehat{G}$ be the natural homomorphism of a group to its profinite completion. The group $G$ is good if, for every finite discrete $\widehat{G}$-module $M$, the homomorphism induced on cohomology $\iota^{\ast}\colon\thinspace H^{i}(\widehat{G},M)\to H^{i}(G,M)$ is an isomorphism for all $i\geq 0$. Let us recall the following definition (cf. [5, Definition 1.2]):
Definition 2.3.
A group $G$ is of finite virtual cohomological type if:
  1. (i)
    $G$ has finite virtual cohomological dimension;
  2. (ii)
    for every finite $G$-module $M$, the cohomology $H^{i}(G,M)$ is finite for $i\geq 0$.
As a corollary of [5, Theorem A], we then have:
Proposition 2.4.
Let $G$ be a good, residually finite group of finite virtual cohomological type and let $\iota\colon\thinspace G\hookrightarrow\widehat{G}$ be the natural monomorphism. Then, for every finite subgroup $H$ of $G$, we have the identity $C_{\widehat{G}}(\iota(H))=\overline{\iota(C_{G}(H))}$.
Proof.
By [5, (iii) of Theorem A], for all primes $p>0$ and every $p$-torsion element $x$ of $G$, there holds $C_{\widehat{G}}(\iota(x))=\overline{\iota(C_{G}(x))}$. The finite group $H$ is generated by a finite subset $\{x_{1},\dots,x_{k}\}$ of $p$-torsion elements (for, possibly, different primes $p>0$). We then have:
$C_{\widehat{G}}(\iota(H))=\bigcap_{i=1}^{k}C_{\widehat{G}}(\iota(x_{i}))=\bigcap_{i=1}^{k}\overline{\iota(C_{G}(x_{i}))}=\overline{\iota(C_{G}(H))}.$
3. Proof of Theorem A
3.1. The case of a free action
The first case to consider for the proof of Theorem A is the one in which $G$ acts freely on $S$, that is to say $R=B=\emptyset$ and so $\Pi=K$ identifies with a normal finite index subgroup of $\Pi_{/G}=\pi_{1}(S_{/G})$.
The congruence subgroup property for ${\Gamma}_{G}(S)$ is equivalent to the statement that the congruence topology on this group, as defined in Section 2.2, is equivalent to the profinite topology. Since, by Lemma 2.2, the subgroup ${\Gamma}_{G}^{(m)}(S)$ of ${\Gamma}_{G}(S)$ is open for the congruence topology, the congruence subgroup property for ${\Gamma}_{G}(S)$ follows if we show that the homomorphism:
$\hat{\rho}_{S,G}^{(m)}\colon\thinspace{\widehat{\Gamma}}_{G}^{(m)}(S)\to\operatorname{Out}({\widehat{\Pi}}),$
induced by the restriction of $\rho_{S,G}$ to ${\Gamma}_{G}^{(m)}(S)$, is injective for some $m\geq 2$.
Since, by hypothesis, the representation $\hat{\rho}_{S_{/G}}\colon\thinspace{\widehat{\Gamma}}(S_{/G})\to\operatorname{Out}({\widehat{\Pi}}_{/G})$ is faithful, its restriction $\hat{\rho}_{S_{/G}}^{(m)}\colon\thinspace{\widehat{\Gamma}}^{K,(m)}\to\operatorname{Out}({\widehat{\Pi}}_{/G})$ is faithful as well for all $m\geq 2$.
For $m\geq 3$, the natural isomorphism ${\Gamma}_{G}^{(m)}(S)\cong{\Gamma}^{K,(m)}$ induces an isomorphism of profinite completions ${\widehat{\Gamma}}_{G}^{(m)}(S)\cong{\widehat{\Gamma}}^{K,(m)}$ and, since ${\widehat{\Pi}}\cong\widehat{K}$ identifies with a normal finite index subgroup of ${\widehat{\Pi}}_{/G}$ and the representations $\hat{\rho}_{S,G}^{(m)}$ and $\hat{\rho}_{S_{/G}}^{(m)}$ are compatible via these isomorphisms, the case of a free action of Theorem A follows from the well known lemma:
Lemma 3.1.
Let ${\widehat{\Pi}}$ be a nonabelian profinite surface group, $N$ an open normal subgroup of ${\widehat{\Pi}}$ and $f\in\operatorname{Aut}({\widehat{\Pi}})$. If $f$ restricts to the identity on $N$, then $f$ is the identity.
Proof.
The proof is essentially the same of [1, Lemma 8]. Let $a_{1},a_{2}\in{\widehat{\Pi}}$ be distinct simple generators. Then, for some $n\in\mathbb{N}^{+}$, we have that $a_{1}^{n},a_{2}^{n}\in N$ and, since $N$ is normal, also that $xa_{i}^{n}x^{-1}\in N$, for all $x\in{\widehat{\Pi}}$ and $i=1,2$.
By hypothesis, we have that $f(xa_{i}^{n}x^{-1})=f(x)a_{i}^{n}f(x)^{-1}=x^{-1}a_{i}^{n}x$ and then $f(x)x^{-1}\in N_{\widehat{\Pi}}(\langle a_{i}^{n}\rangle)$, for $i=1,2$. Since, by [4, Proposition 3.6], there holds $N_{\widehat{\Pi}}(\langle a_{i}^{n}\rangle)=\langle a_{i}\rangle$, for $i=1,2$, and $\langle a_{1}\rangle\cap\langle a_{2}\rangle=\{1\}$, the conclusion follows. ∎
3.2. The general case
The completions of the groups ${\Gamma}(S)$ and ${\Gamma}_{G}(S)$ with respect to the congruence topologies which we introduced in Section 2.2 are called procongruence completions and denoted by ${\check{\Gamma}}(S)$ and ${\check{\Gamma}}_{G}(S)$, respectively. The group ${\check{\Gamma}}(S)$ is also called the procongruence mapping class group. The congruence subgroup problems for ${\Gamma}(S)$ and ${\Gamma}_{G}(S)$ can then be formulated as the questions of whether the natural epimorphisms ${\widehat{\Gamma}}(S)\to{\check{\Gamma}}(S)$ and ${\widehat{\Gamma}}_{G}(S)\to{\check{\Gamma}}_{G}(S)$ are, respectively, isomorphisms.
By the results of Section 3.1 and with the notations of Section 1, we know that there is a natural isomorphism ${\widehat{\Gamma}}_{G}(S^{\circ})\cong{\check{\Gamma}}_{G}(S^{\circ})$. Therefore, in order to prove the general case of Theorem A, it is enough to show that the natural isomorphism $q^{\circ}\colon\thinspace{\Gamma}_{G}(S^{\circ})\stackrel{{\scriptstyle\sim}}{{\to}}{\Gamma}_{G}(S)$, induced by the natural epimorphism $q^{\circ}\colon\thinspace{\Gamma}(S^{\circ})\to{\Gamma}(S)$ (cf. Proposition 2.1), induces an isomorphism between the procongruence completions: $\check{q}^{\circ}\colon\thinspace{\check{\Gamma}}_{G}(S^{\circ})\stackrel{{\scriptstyle\sim}}{{\to}}{\check{\Gamma}}_{G}(S)$.
The set $R$ of ramification points of the quotient map $p_{G}\colon\thinspace S\to S_{/G}$ determines a point, which we also denote by $R$, on the unordered configuration space $S[r]$ of $r$ distinct points on the surface $S$, where $r:=\sharp R$.
Let $\Pi[r]:=\pi_{1}(S[r],R)$. From [6, Corollary 4.7 and Proposition 4.9], it then follows that there is a short exact sequence of profinite groups:
(4) $1\to{\widehat{\Pi}}[r]\to{\check{\Gamma}}(S^{\circ})\stackrel{{\scriptstyle\check{q}^{\circ}}}{{\to}}{\check{\Gamma}}(S)\to 1$
In order to prove Theorem A, it is then enough to show that the subgroup ${\check{\Gamma}}_{G}(S^{\circ})$ of ${\check{\Gamma}}(S^{\circ})$ has trivial intersection with the image of ${\widehat{\Pi}}[r]$ in ${\check{\Gamma}}(S^{\circ})$. Since ${\check{\Gamma}}_{G}(S^{\circ})$ is contained in the centralizer $Z_{{\check{\Gamma}}(S^{\circ})}(G)$, the conclusion follows if we prove that $Z_{{\check{\Gamma}}(S^{\circ})}(G)\cap{\widehat{\Pi}}[r]=\{1\}$.
By Proposition 2.1 and the short exact sequence (4), the subgroup $G$ of ${\Gamma}(S^{\circ})$ has trivial intersection with ${\widehat{\Pi}}[r]$, so that ${\widehat{\Pi}}[r]\cdot G\cong{\widehat{\Pi}}[r]\rtimes G$ identifies with the profinite completion of $\Pi[r]\cdot G\cong\Pi[r]\rtimes G$. We are then reduced to prove the following lemma:
Lemma 3.2.
There holds $Z_{{\widehat{\Pi}}[r]\cdot G}(G)=Z(G)$.
Proof.
Proposition 2.1 implies that ${\Gamma}_{G}(S^{\circ})\cap\Pi[r]=\{1\}$ and then that $Z_{\Pi[r]\cdot G}(G)=Z(G)$. Since $\Pi[r]\cdot G$ contains, as a finite index subgroup, the fundamental group of the ordered configuration space of $n$ points on the surface $S$, which is a good, residually finite group of finite virtual cohomological type, by the Lyndon-Hochschild-Serre spectral sequence, it follows that $\Pi[r]\cdot G$ has the same properties. Hence, we conclude by Proposition 2.4. ∎
4. The congruence subgroup property for relative mapping class groups
Let us observe that, for $\partial S\neq\emptyset$, the natural representation ${\Gamma}(S,\partial S)\to\operatorname{Out}(\Pi)$ is not faithful. Hence, the congruence subgroup property, as formulated in Section 1 does not make sense for relative mapping class groups. For this reason, in [7, Section 2.7], it was introduced the following construction.
Let ${\sigma}:=\{\delta_{1},\ldots,\delta_{k}\}$ be the set of connected components of $\partial S$ and let $\widetilde{S}$ be the surface (without boundary) obtained from $S$ attaching to each curve $\delta_{i}$ a $2$-punctured closed disc $D_{i}$, for $i=1,\dots k$. Let then ${\Gamma}(\widetilde{S})_{\sigma}$ be the stabilizer in ${\Gamma}(\widetilde{S})$ of the set of isotopy classes of the curves contained in ${\sigma}$. The relative mapping class group ${\Gamma}(S,\partial S)$ acts on the set ${\sigma}$ and the associated representation in the symmetric group ${\Gamma}(S,\partial S)\to\Sigma_{k}$ is surjective. There is then a natural isomorphism:
(5) ${\Gamma}(\widetilde{S})_{\sigma}\cong(\Sigma_{2})^{k}\rtimes{\Gamma}(S,\partial S),$
where the mapping class group ${\Gamma}(S,\partial S)$ acts on the product of symmetric groups $(\Sigma_{2})^{k}$ by permuting its factors through the epimorphism ${\Gamma}(S,\partial S)\to\Sigma_{k}$ defined above.
The isomorphism (5) identifies the relative mapping class group ${\Gamma}(S,\partial S)$ with a subgroup of the mapping class group ${\Gamma}(\widetilde{S})$. We then define the congruence topology on ${\Gamma}(S,\partial S)$ as the topology induced, via this embedding, by the congruence topology of ${\Gamma}(\widetilde{S})$. The relative procongruence mapping class group ${\check{\Gamma}}(S,\partial S)$ is the associated completion of ${\Gamma}(S,\partial S)$.
The congruence subgroup problem for ${\Gamma}(S,\partial S)$ then asks whether the natural epimorphism ${\widehat{\Gamma}}(S,\partial S)\to{\check{\Gamma}}(S,\partial S)$ is an isomorphism.
For a finite group acting faithfully on $S$, we let ${\Gamma}_{G}(S,\partial S)$ be the centralizer of $G$ in ${\Gamma}(S,\partial S)$. We define the congruence topology on ${\Gamma}_{G}(S,\partial S)$ as the one induced by the congruence topology of ${\Gamma}(S,\partial S)$ and we denote by ${\check{\Gamma}}_{G}(S,\partial S)$ the associated completion. We can then formulate the corresponding congruence subgroup problem for ${\Gamma}_{G}(S,\partial S)$.
There is a natural short exact sequence:
(6) $1\to\prod_{i=1}^{k}\tau_{{\delta}_{i}}^{\mathbb{Z}}\to{\Gamma}(S,\partial S)\to{\Gamma}(S)\to 1,$
where $\tau_{{\delta}_{i}}^{\mathbb{Z}}\cong\mathbb{Z}$ is the cyclic subgroup of ${\Gamma}(S,\partial S)$ topologically generated by the Dehn twist about the curve ${\delta}_{i}$, for $i=1,\dots k$. In [7, Section 2.7], it was then observed that the procongruence completion of ${\Gamma}(S,\partial S)$ induces the short exact sequence:
(7) $1\to\prod_{i=1}^{k}\tau_{{\delta}_{i}}^{\widehat{\mathbb{Z}}}\to{\check{\Gamma}}(S,\partial S)\to{\check{\Gamma}}(S)\to 1,$
where $\tau_{{\delta}_{i}}^{\widehat{\mathbb{Z}}}\cong{\widehat{\mathbb{Z}}}$ is the closure of $\tau_{{\delta}_{i}}^{\mathbb{Z}}$ in the procongruence completion ${\check{\Gamma}}(S,\partial S)$, for $i=1,\dots k$.
From the above two exact sequences, it follows, in particular, that the congruence subgroup property for the mapping class group ${\Gamma}(S)$ implies the same property for the relative mapping class group ${\Gamma}(S,\partial S)$. A similar argument yields the following generalization of Theorem A and Corollary B:
Theorem 4.1.
Let $S$ be an oriented connected hyperbolic surface of finite type and $G$ a finite group acting faithfully on $S$. If the congruence subgroup property holds for the mapping class group ${\Gamma}(S_{/G}^{\circ})$ of the surface $S_{/G}^{\circ}$, then it holds for the centralizer ${\Gamma}_{G}(S,\partial S)$ of $G$ in the relative mapping class group ${\Gamma}(S,\partial S)$. In particular, for $g(S_{/G})\leq 2$, there is a natural isomorphism ${\widehat{\Gamma}}_{G}(S,\partial S)\cong{\check{\Gamma}}_{G}(S,\partial S)$.
Proof.
The group $G$ acts by conjugation on the normal abelian subgroup $\prod_{i=1}^{k}\tau_{{\delta}_{i}}^{\mathbb{Z}}$ of ${\Gamma}(S,\partial S)$. Let us denote by $A^{G}$ the invariants for the action of a group $G$ on an abelian group $A$. We then have the short exact sequence:
$1\to(\prod_{i=1}^{k}\tau_{{\delta}_{i}}^{\mathbb{Z}})^{G}\to{\Gamma}_{G}(S,\partial S)\to{\Gamma}_{G}(S)\to 1$
Similarly, for the procongruence completion, we get the short exact sequence:
$1\to(\prod_{i=1}^{k}\tau_{{\delta}_{i}}^{\widehat{\mathbb{Z}}})^{G}\to{\check{\Gamma}}_{G}(S,\partial S)\to{\check{\Gamma}}_{G}(S)\to 1,$
where the subgroup of invariants $(\prod_{i=1}^{k}\tau_{{\delta}_{i}}^{\widehat{\mathbb{Z}}})^{G}$ identifies with the closure of the subgroup $(\prod_{i=1}^{k}\tau_{{\delta}_{i}}^{\mathbb{Z}})^{G}$ in ${\check{\Gamma}}_{G}(S,\partial S)$. Theorem 4.1 then follows from the two above exact sequences and Theorem A if we show that the profinite group $(\prod_{i=1}^{k}\tau_{{\delta}_{i}}^{\widehat{\mathbb{Z}}})^{G}$ is the profinite completion of the group $(\prod_{i=1}^{k}\tau_{{\delta}_{i}}^{\mathbb{Z}})^{G}$. But this is clear since the profinite completion of a finitely generated abelian group induces the profinite completion on each of its subgroups. ∎
5. Proof of Theorem C
By [5, Corollary C], it is enough to prove that, for $g(S)\leq 2$, the relative mapping class group ${\Gamma}(S,\partial S)$ is good and that, for every finite index subgroup $G$ of ${\Gamma}(S,\partial S)$, the following conditions are satisfied:
  1. (i)
    the normalizer $N_{{\Gamma}(S,\partial S)}(G)$ is also a good group;
  2. (ii)
    the closure of the normalizer $N_{{\Gamma}(S,\partial S)}(G)$ in ${\widehat{\Gamma}}(S,\partial S)$ coincides with the profinite completion of $N_{{\Gamma}(S,\partial S)}(G)$.
Since ${\Gamma}(S)$ is a good group for $g(S)\leq 2$, from Lyndon-Hochschild-Serre spectral sequence and the short exact sequences (6) and (7), we deduce that the relative mapping class group ${\Gamma}(S,\partial S)$ is also good for $g(S)\leq 2$.
There is a version of the short exact sequence (2) for relative mapping class groups:
$1\to G\to N_{{\Gamma}(S,\partial S)}(G)\to{\Gamma}(S_{/G}^{\circ},\partial S_{/G}^{\circ})_{K}\to 1$
By the previous remark, this implies that $N_{{\Gamma}(S,\partial S)}(G)$ is a good group for $g(S_{/G})\leq 2$. Theorem 4.1 then implies in particular that, for $g(S_{/G})\leq 2$, the subspace topology on $N_{{\Gamma}(S,\partial S)}(G)\subseteq{\Gamma}(S,\partial S)$, induced by the profinite topology of ${\Gamma}(S,\partial S)$, is the full profinite topology on this group.